On a radar screen, a plane located at A(-2,4) is flying toward B(4,3) Another plane, located at C(-3,1), is flying toward D(3,0). Are the planes' paths perpendicular? Explain.
1
Expert's answer
2019-12-11T14:15:05-0500

Let's consider the screen of a radar as Descartes's coordinate plane (coordinates of points A, B, C, D are given). Further, points A(хA , yA), B(хB , yB),C(хC , yC), D(хD , yD) will look like


A(-2; 4);B(4; 3);C(-3;1);D(3;0);


(1 plane) yA = kAB*xA + lAB=> lAB =- yA + kAB*xA => yA = kAB*xA + yB - kAB*xBkAB = (yA - yB) / (xA - xB) => kAB=(4-3)/(-2-4)=1/-6=-1/6
(2 plane) yC = kCD*xC + lCD => yD = kCD*xD + lCDlCD = - yD + kCD*xD =>yC = kCD*xC + yD - kCD*xDkCD = (yC - yD) / (xC - xD) => kCD=(1-0)/(-3-3)=-1/6The slopes of plane's lines are equal

kAB=kCD => their paths are parallel